# A Survey on Vision-Language-Action Models for Embodied AI

### v2

@@ IV Datasets and Benchmarks
EgoPlan [264] 3,355 2,432 + 923 Epic-Kitchen + Ego4D Multiple-choice (1 out of 4) No GPT-4 filtering, question template, human verification Acc
@@
EgoCOT [101] 5,246 plans 129 Ego4D Caption and planning (not QA) No ChatGPT PPL
LoTa-Bench [267] 17K instruction-plan pairs AI2-THOR, VirtualHome Goal condition - ALFRED + WAH-NL Success
Furthermore, some datasets and benchmarks may not be directly dedicated to robot manipulation and navigation, but they are aimed at other relevant abilities for embodied AI, such as spatial reasoning, physics understanding, as well as world knowledge. One of the most prominent embodied tasks is embodied question answering (EQA), as summarized in Table VII. EQA is akin to previous visual question answering and video question answering tasks but differs in that the agent can actively explore the environment before providing an answer. EmbodiedQA [258] and IQUAD [259] were among the first works to introduce this task. MT-EQA [260] focuses on questions involving multiple targets, increasing the complexity of understanding and answering the questions. MP3D-EQA [261] converts previous RGB input to point clouds, testing the 3D perception capability. However, active exploration requires access to a simulator, limiting the types of data that can be used, such as real-world videos. Therefore, some EQA benchmarks do not involve active exploration. EgoVQA [262] shifts the focus of VQA to egocentric videos. EgoTaskQA [263] emphasizes spatial, temporal, and causal relationship reasoning. EQA-MX [266] focuses on multimodal expressions (MX), including regular verbal utterances and nonverbal gestures like eye gaze and pointing. OpenEQA [265] evaluates seven main categories, including functional reasoning and world knowledge, which were not covered in previous benchmarks. EgoPlan-Bench [264] and EgoCOT [101] measure the model’s ability to generate task plans, using metrics like accuracy and perlexity. PlanBench [268, 269] comprehensively assesses various aspects of task planning ability, such as cost optimality, plan verification, and replanning, etc. LoTa-Bench [267] directly evaluates task planning by executing the generated plans in simulators and calculating the success rate.
V Challenges and Future Directions
Vision-language-action (VLA) models face several persistent challenges in the domain of Robotics, necessitating focused attention and concerted research efforts:
Scarcity of Robotic Data. Obtaining sufficient real-world robotic data remains a significant obstacle. Collecting such data is time-consuming and resource-intensive, while relying solely on simulation data exacerbates the sim-to-real gap problem. Diverse real-world robotic datasets require close and extensive collaboration among different institutions. Simulated data depends on the development of more realistic and efficient simulators.
Motion Planning. Current motion planning modules often lack the necessary dexterity to address the complexities in various environments. This limitation hampers robots’ abilities to interact effectively with tools, navigate complex environments, and execute high-precision operations, etc. Overcoming these challenges necessitates the development of more robust motion planning algorithms.
Real-Time Responsiveness. Many robotic applications require real-time decision-making and action execution to meet operational requirements. VLA models should be designed to operate responsively, with minimal latency. Furthermore, the entire stack of a robot system necessitates global optimization, from the high-level task planner down to the motion planner.
Integration of Multiple Modalities. VLAs must process and integrate information from multiple modalities, including vision, language, and action. While significant progress has been made in this regard, achieving optimal integration of these modalities remains an ongoing challenge. Addressing this challenge requires advancements in multi-modal representation learning, fusion techniques, and task-specific adaptation. Expanding beyond mere visual and linguistic capacities, robots can greatly benefit from incorporating modalities like audio or speech. Embracing a broader range of perceptual and communicative abilities empowers robots to engage in more effective collaboration with users.
Generalization to Unseen Scenarios. A truly versatile robotics system should be capable of comprehending and executing natural language instructions across diverse and unseen scenarios. Achieving this level of generalization, akin to the capabilities demonstrated by ChatGPT in natural language processing, requires robustness to variations in instructions, environments, objects, and embodiments. This necessitates the development of adaptable and scalable VLA architectures.
Long-Horizon Task Execution. Single instructions can often translate into long-horizon tasks for robots, such as the instruction “clean the room” which involves object rearrangement, floor sweeping, table wiping, and more. Successfully executing such tasks requires the robot to plan and execute a sequence of low-level actions over extended time horizons. While current high-level task planners have achieved initial success, they still fall short in many scenarios because most LLMs are not tuned to embodied tasks. Addressing this challenge entails devising efficient planners that possess powerful perception abilities and a wide range of common sense.
Foundation Model. The exploration of foundation models for VLAs in robot tasks remains uncharted territory, primarily due to the diverse embodiments, environments, and tasks encountered in robotics. Isolated datasets and evaluation setups further compound this challenge. To establish a robust foundation VLA model for robotics, it is imperative to leverage internet-scale embodied datasets and state-of-the-art multi-modal models.
Multi-agent Systems. The large-scale deployment of robots is highly desirable in many scenarios, but it introduces the challenge of multi-agent cooperation. Effective communication, task distribution among multiple robots can be difficult [270]. Fleet heterogeneity may further complicate these issues. While multi-agent systems offer benefits like distributed perception, collaborative fault recovery, and autonomous fleet maintenance, realizing the full potential of these advantages remains elusive.
Benchmarks. Although numerous benchmarks exist for evaluating low-level control policy VLAs, they often vary significantly in terms of the skills they assess. Additionally, the objects and scenes included in these benchmarks are typically limited by what the simulator can provide. To more comprehensively evaluate VLA models, there is a need for benchmarks that cover a diverse set of skills based on realistic simulators. For high-level task planner VLAs, many benchmarks claim to measure planning capability, often in the form of question-answering tasks. However, it is more desirable to evaluate the high-level task planner together with the low-level control policy to execute long-horizon tasks and measure success rate, rather than relying on isolated measures of the planner alone. This approach provides a more holistic assessment of the VLA system’s capabilities.
@@ V Challenges and Future Directions
Safety Considerations. Safety is paramount in robotics, as robots interact directly with the real world. Ensuring the safety of robotic systems requires the integration of real-world commonsense and complex reasoning into their development and deployment processes. This involves the incorporation of robust safety mechanisms, risk assessment frameworks, and human-robot interaction protocols. Interpretability and expandability of VLA decision-making processes are also crucial for enhancing robot safety through error diagnosis and troubleshooting.
@@
Ethical and Societal Implications. The deployment of robots has always raised various ethical, legal, and societal concerns. These include risks related to privacy, safety, job displacement, bias in decision-making, and the impact on social norms and human relationships. Effective regulation plays a crucial role in promoting the ethical use of robots.
VI Conclusion
VLA policy holds immense promise for enabling embodied AI to interact effectively with the world around them. Recent advances have demonstrated the capabilities of these models in completing complex tasks under diverse conditions. However, significant challenges remain in terms of generalization, efficiency, and safety. Further research is needed to address these challenges and pave the way for the widespread adoption of VLA-powered robots in real-world applications.

### v4

@@ V-F Embodied Question Answering Benchmarks
VI Challenges and Future Directions
Safety first
@@ VI Challenges and Future Directions
Safety is paramount in robotics, as robots interact directly with the physical world. Ensuring the safety of robotic systems requires the integration of real-world commonsense. This involves the incorporation of robust safety guardrails, risk assessment frameworks, and human-robot interaction protocols. RLHF and “evaluation without execution” can also significantly lower safety risks [37]. Interpretability and expandability of VLA decision-making processes are also crucial for enhancing robot safety through error diagnosis and troubleshooting.
Datasets & Benchmarks
In addition to the issues discussed in §V, comprehensive benchmarks that cover a wide range of skills, objects, embodiments, and environments remain to be developed. Moreover, metrics beyond the success rate are needed for a fine-grained diagnosis of issues in VLA models, as highlighted by EAI [222] for LLMs.
Foundation Models & Generalization
VLA foundation models or robotic foundation models (RFM) for embodied AI remains an open research topic primarily due to the diversity in embodiments, environments, and tasks. Many have made significant progress [139, 147], but still lacks generalization capability on par with LLMs in NLP. Attaining such a level of generalization is very challenging, as it requires the development of many core AGI capabilities.
Multi-Modality
VLAs inherit many challenges associated with multimodal models, such as obtaining useful embeddings and aligning different modalities. Current approaches, like ImageBind [231] and LanguageBind [232], align different modalities to the image or language embedding space, respectively. Within a unified embedding space, MLLMs can accommodate diverse modalities and become more general. However, whether focusing on embeddings alone is sufficient remains under debate. Although beyond the scope of VLAs, other modalities—such as audio [60] and haptics [233]—have proven useful for certain embodied AI applications. Incorporating additional modalities is often desirable, but they also introduce extra complexity into the model design.
Framework for Long-Horizon Tasks
The hierarchical framework is currently the most practical approach for long-horizon tasks. However, it increases system complexity and potential points of failure. Frequent task execution failures can trigger re-planning, which can cause significant latency. Moreover, monolithic task planners share similar architectures with LVLAs, so employing two large models can be redundant and may hinder scalability. Additionally, although modular task planners typically do not require training, they are not plug-and-play: language-based models may generate subtasks that control policies cannot execute, whereas code-based models require modules to be pre-wrapped in APIs manually. Therefore, developing a unified framework that directly translates long-horizon tasks into low-level control signals in an end-to-end fashion is worth exploring.
Real-Time Responsiveness
Unlike conversational AI, many robotic applications require real-time decision-making to respond to dynamic environments. If inference time cannot keep pace with environmental changes, the model may generate obsolete actions repeatedly. However, current VLA models—especially LVLAs and task planners—face a trade-off between speed and capacity. Novel mechanisms are thus needed to strike an optimal balance.
Multi-agent Systems
Cooperative multi-agent systems offer benefits such as distributed perception and collaborative fault recovery. However, they also face challenges, including effective communication, coordinated dispatching, and fleet heterogeneity. In certain scenarios, individual agents may have conflicting goals, which further increases complexity.
Ethical and Societal Implications
@@
Robotics has always raised various ethical, legal, and societal concerns. These include risks related to privacy, safety, job displacement, bias in decision-making, and the impact on social norms and human relationships.
Applications
Most current VLAs focus on household and industrial settings [2], which shape the available datasets and models. Given the potential of VLA models, a wide range of new applications is possible. One particularly important application is in healthcare, such as surgical robots [234] and care robots [235]. Healthcare demands higher safety and privacy standards and may necessitate novel techniques such as human-in-the-loop (HITL) control and federated learning. Moreover, due to the significant domain gap, specialized vision models might be needed for medical images.
VII Conclusion
Vision-language-action models hold immense promise for enabling embodied agents to interact with the physical world and fulfill users’ instructions. This paper is the first survey to review large VLAs alongside other generalized VLAs. Our taxonomy provides a high-level overview of three main lines of research: key components, control policies, and task planners. We meticulously summarize and compare their model designs and technical details. Additionally, we highlight essential resources for training and evaluating VLAs, such as datasets, simulators, and benchmarks. We hope this survey captures the rapidly evolving landscape of embodied AI and inspires future research.